# Multivariate Causal Effects: a Bayesian Causal Regression Factor Model

### v2

@@ Abstract.
footnotetext: dafne__\__zorzetto@brown.edu
@@ 1. Introduction
Wildfires have emerged as a critical environmental and public health concern, significantly deteriorating air quality through the emission of smoke and associated pollutants (Childs et al., 2022; Krasovich Southworth et al., 2025). In the United States, wildfire smoke exposure has increased significantly, with the burned area quadrupling over the past four decades, a trend expected to continue as climate conditions become increasingly favorable to fire activity. Wildfire smoke can rise high into the atmosphere and travel long distances, affecting air pollution levels and health risks far from the source (Burke et al., 2021; O’Dell et al., 2021). As a result, wildfire pollution has received considerable attention in environmental health research (Reisen et al., 2015; Burke et al., 2021; Krasovich Southworth et al., 2025; Aguilera et al., 2021).
@@
A growing body of literature has examined the effects of wildfire smoke on particulate matter (pm2.5), highlighting the need to quantify the proportion of pm2.5 attributable to wildfires (Burke et al., 2023; Childs et al., 2022), given the well-documented adverse health effects of elevated pm2.5 exposure (Carone et al., 2020; U.S. Environmental Protection Agency, 2022). However, wildfire smoke is a complex mixture of pollutants that undergoes physical and chemical transformations during atmospheric transport (Reisen et al., 2015). Understanding how wildfire smoke influences the concentration levels of various chemical species in pm2.5is therefore a pressing challenge.
@@
While previous studies have focused on the impact of wildfire smoke on individual chemical components —such as phosphorus and nitrogen (Spencer and Hauer, 1991), lead (Odigie and Flegal, 2014), bromine and other metals (Li et al., 2023)— the recent work of Krasovich Southworth et al. (2025) has provided a more comprehensive analysis by quantifying the contribution of wildfire smoke to the concentration of 27 different chemical species in pm2.5 using a standard linear regression model. Their analysis focuses primarily on accounting for temporal variation and estimating the contribution of wildfire smoke to each chemical species in pm2.5. However, it does not jointly model chemical species, limiting information sharing across species, and does not adopt a causal inference perspective, restricting its ability to disentangle the causal effect of wildfire smoke.
@@
A causal inference framework is crucially needed to rigorously assess and quantify the effect of wildfire smoke on changes in chemical concentration levels, providing a deeper understanding of the associations highlighted in previous studies. Moreover, incorporating the high correlation structure among the different chemical species is critical to capturing deeper insights into the composition of smoke and its broader environmental effects.
@@
Traditional Bayesian factor models have focused primarily on priors for factor loadings (Bhattacharya and Dunson, 2011; Schiavon et al., 2022; Frühwirth-Schnatter et al., 2024; De Vito et al., 2021), typically assuming standard normal distributions for factor scores. Recent works (Zorzetto et al., 2024; Bortolato and Canale, 2024) have begun exploring priors on factor scores, since they define the subject-specific contribution of each factor to the variation, making them essential in many applications. However, introducing a causal inference framework in Bayesian factor models poses new challenges due to the fundamental missing data problem of causal inference (Rubin, 1986). Our proposed prior addresses this by allowing a covariate-dependent structure in treatment-specific scores, enabling better estimation of latent confounding patterns. In particular, our prior leverages the flexibility of the Dependent Dirichlet Process (DDP) (Quintana et al., 2022), building on a growing body of work in causal inference that employs DDP priors (e.g., Roy et al., 2024; Hu et al., 2023; Zorzetto et al., 2024), though primarily in the context of scalar outcomes.
The paper is organized as follows. Section 2 introduces the causal inference setup, including the estimands of interest, causal assumptions, and identification strategy. Section 3 introduces our Bayesian factor model for causal inference, highlighting the innovation in our prior specification for factor scores and its critical role in this context. In Section 4, we assess the performance of the model performance through simulations, comparing it to state-of-art approaches in causal inference and evaluating the importance of our proposed prior for the factor scores, compared to the factor model with the standard prior. In Section 5, we apply our approach to estimate the causal effect of the presence of wildfire smoke on the concentration of 27 chemical species in pm2.5 across the United States, using data from common sources of Krasovich Southworth et al. (2025): F.E.D. IMPROVE (2024) and Childs et al. (2022). We conclude the paper with a discussion and future directions in Section 6.
@@ 2. Causal setup
Let $i$ denote the study unit, with $i=1,\dots,n$. Foe each unit, let $T_{i}\in\{0,1\}$ the binary treatment random variable with observed value $t_{i}$, let ${\bf X}_{i}=(X_{1i},\dots,X_{pi})^{T}$ be the vector of $p$ observed covariates (confounders), which may be either continuous or discrete, and let $\mathbf{Y}_{i}\in\mathbb{R}^{q}$ be the multivariate outcome of dimension $q$. The outcome components are potentially highly correlated, a structure acknowledged by our Bayesian causal factor model. Throughout the paper, bold letters denote vectors.
Following the definition of the Rubin Causal Model (Rubin, 1974), we make the Stable Unit Treatment Value Assumption (SUTVA; Rubin, 1980), which is a combination of two assumptions: no interference among the units, i.e. the potential outcome values from unit $i$ do not depend on the treatment applied to other units, and consistency, i.e. each observed outcome is the individual’s potential outcome under their treatment. Under SUTVA, we define the potential outcomes, for unit $i$, as $\{{\bf Y}_{i}(0),{\bf Y}_{i}(1)\}$, where ${\bf Y}_{i}(t)=(Y_{i1}(t),\dots,Y_{iq}(t))^{T}$ for $t\in\{0,1\}$. Specifically, ${\bf Y}_{i}(0)\in{\mathds{R}}^{q}$ is the $q$-variate outcome when the unit $i$ is assigned to the control group and ${\bf Y}_{i}(1)\in{\mathds{R}}^{q}$ is the $q$-variate outcome when the unit $i$ is assigned to the treatment group. Under consistency, therefore the observed outcome can be written as:
${\bf Y}_{i}=(1-T_{i})\,{\bf Y}_{i}(0)+T_{i}\,{\bf Y}_{i}(1).$
In practice, for $i=1,\dots,n$, we observe only ${\bf y}_{i}\in{\mathds{R}}^{q}$, the realization of the random variable ${\bf Y}_{i}=(Y_{i1},\dots,Y_{iq})$. Conversely, we do not observe the missing outcome ${\bf Y}_{i}^{mis}\in{\mathds{R}}^{q}$ defined as ${\bf Y}_{i}^{mis}=T_{i}\,{\bf Y}_{i}(0)+(1-T_{i})\,{\bf Y}_{i}(1)$.
To concisely define the correlated nature of the multidimensional outcome, for each unit $i$, we assume an unmeasured variable ${U}_{i}$ that affects the treatment $T_{i}$ and the potential outcome $\{{\bf Y}_{i}(0),{\bf Y}_{i}(1)\}$. Under the assumption of our proposed Bayesian causal factor model, formally introduced in Section 3. Bayesian Causal Factor Model, we model both factual and counterfactual latent factors $\{{\bf L}_{i0},{\bf L}_{i1}\}$ that capture the indirect effect of $U_{i}$ on the potential outcomes $\{{\bf Y}_{i}(0),{\bf Y}_{i}(1)\}$, as illustrated in Figure 1. Figure 1 shows the three versions of the resulting conditional independence structure considered in our model.
@@
Positivity. Each unit $i\in\{1,\dots,n\}$ has a non-zero probability of receiving each treatment level,
$10<\Pr(T_{i}=1\mid{\bf X}_{i}={\bf x}_{i})<$ 1 .
Conditional ignorability. The assignment of treatment is random in each group of units characterized by observed covariates ${\bf x}$ and latent factors scores ${\bf l}_{0}$ and ${\bf l}_{1}$, such that for each unit $i\in\{1,\dots,n\}$
$\{{\bf Y}_{i}(1),{\bf Y}_{i}(0)\}\perp T_{i}\mid\{{\bf X}_{i}={\bf x}_{i},{\bf L}_{i0}={\bf l}_{i0},{\bf L}_{i1}={\bf l}_{i1}\}.$
Indirectness of unmeasured confounding effects. The unmeasured variable $U$ affects the potential outcomes only indirectly, through the latent factor and the observed confounders, such as
$\mathds{P}({\bf Y}(1),{\bf Y}(0)\mid{\bf X},{\bf L}_{0},{\bf L}_{1},U)=\mathds{P}({\bf Y}(1),{\bf Y}(0)\mid{\bf X},{\bf L}_{0},{\bf L}_{1})$
This assumption can be visualized in Figure 1 as the absence of a direct arrow between the $U$ and ${\bf Y}$.
@@
${\mathds{E}}[{\bf Y}(1)-{\bf Y}(0)]$
$\quad=\int_{\mathcal{X}}\int_{\mathbb{R}^{J_{0}}}\int_{\mathbb{R}^{J_{1}}}\bigg{(}{\mathds{E}}[{\bf Y}\mid T=1,{\bf X}={\bf x},{\bf L}_{0}={\bf l}_{0},{\bf L}_{1}={\bf l}_{1}]$
$\quad\quad\quad\quad-{\mathds{E}}[{\bf Y}\mid T=0,{\bf X}={\bf x},{\bf L}_{0}={\bf l}_{0},{\bf L}_{1}={\bf l}_{1}]\bigg{)}$
×$\quad\quad\quad\quad\times\mathds{P}({\bf L}_{0}={\bf l}_{0},{\bf L}_{1}={\bf l}_{1}\mid{\bf X}={\bf x})\mathds{P}({\bf X}={\bf x})d{\bf l}_{0},d{\bf l}_{1}d_{\bf x};$
@@
${\mathds{E}}[{\bf Y}\mid t,x,l_{0},l_{1}]$
[…]
@@ 3. Bayesian Causal Factor Model
To address the multivariate outcome, we propose an approach that adopt a factor regression model (Carvalho et al., 2008; Avalos-Pacheco et al., 2022; De Vito and Avalos-Pacheco, 2023), which uses factor analysis to model the potential outcomes given the treatment and the confounders. Rather than performing multiple univariate analyses, our factor analysis approach (i) improves statistical power by leveraging the correlation structure in the outcome, and (ii) facilitates the recovery of information from the unmeasured variable ${U}$, under the three cases considered in Figure 1.
@@
(2) $g_{t}({\bf x}_{i},{\bf l}_{it})=\boldsymbol{\mu}_{t}+{\bf B}_{t}{\bf x}_{i}+\Lambda_{t}{\bf l}_{it}+\boldsymbol{\xi}_{it},$
$\boldsymbol{\xi}_{it}\sim\mathcal{N}_{q}(0,\Psi_{t}).$
The outcome ${\bf Y}_{i}$ depends on treatment $t$, covariates ${\bf x}_{i}$, and treatment-specific latent factors ${\bf l}_{it}\in\mathbb{R}^{J_{t}}$, where $J_{t}$ denotes the number of factors for treatment $t$ (assumed unknown). The matrix $\Lambda_{t}\in\mathbb{R}^{q\times J_{t}}$ represents treatment-specific factor loadings, and $\Psi_{t}\in\mathbb{R}^{q\times q}$ is assumed to be a diagonal matrix with elements $(\psi_{1t},\dots,\psi_{qt})$ as in standard factor analysis literature (Thurstone, 1931; Anderson-Cook, 2004).
@@
where $\boldsymbol{\beta}_{tj}$ denotes the $j$-th row of the matrix ${\bf B}_{t}$ in (2), such that $\boldsymbol{\beta}_{tj}=\{\beta_{tj1},\dots,\beta_{tjp}\}$ and $\beta_{tjk}$ is the element $({\bf B}_{t})_{j,k}$ with $k\in\{1,\dots,p\}$. We assume that $\Sigma_{m_{t}}$ and $\Sigma_{\beta_{t}}$ are diagonal matrices for each $t$.
@@
The flexibility of the DDP has made it widely adopted in causal inference settings involving scalar outcomes. For example, Zorzetto et al. (2024) adopt the probit stick-breaking process (Rodriguez and Dunson, 2011) to identifies subpopulations that characterize the heterogeneity in causal effects. Roy et al. (2018) and Roy et al. (2024) leverage the enriched Dirichlet process (Wade et al., 2011) estimate causal effects in the presence of missing-at-random covariates and for mediation analysis, respectively. In addition, Schwartz et al. (2011) and Zorzetto et al. (2024) specify a DDP prior within the principal stratification framework.
In our factor model, our DDP prior has two characteristics: (i) it can written as infinite mixture of Gaussian distributions, and (ii) the probability of assignment to each mixture component depends on the covariates $X$—called cluster allocation in the DDP literature. Moreover, we specify a DDP prior that allows us to capture heterogeneity at both the factor and treatment levels. In fact, treatment-specific factor scores quantify the score/importance that each unit expresses on the corresponding factor under the specific treatment level. In the environmental application discussed in Section 5. Environmental Application, each factor score of a specific chemical indicates its contribution to explain the correlation among chemicals within the corresponding latent factor (reflecting underlying sources of metal and organic emissions). Notably, each factor score can vary depending on the presence or absence of wildfire smoke, reflecting how wildfire exposure alters the chemical composition of the air.
@@
$\{l_{itj}\mid{\bf x}_{i},t\}\sim h_{tj}(\cdot\mid{\bf x}_{i}),$
(3) $h_{tj}(\cdot\mid{\bf x}_{i})=\int_{\Omega}{\mathcal{H}}(\cdot;\omega)dG_{{\bf x}_{i}}^{(tj)}(\omega),$
$G_{{\bf x}_{i}}^{(tj)}\sim\Pi_{{\bf x}_{i}}^{(tj)},$
for $j\in\{1,\dots,J_{t}\}$ of units $i\in 1,\dots.n\}$, for each $t\in\{0,1\}$. The random probability measure $G_{{\bf x}_{i}}^{(tj)}$ depends on confounders ${\bf x}_{i}$ and is specific to a treatment level $t$ and a factor $j$. Assuming its prior $\Pi_{{\bf x}_{i}}^{(tj)}$ where $\Pi_{{\bf x}_{i}}^{(tj)}$ is a DDP (Quintana et al., 2022), we define a dependent infinite mixture distribution as the probability density distribution of the factors.
@@
Following a single-atom DDP (Quintana et al., 2022) characterization of the random measure $G_{{\bf x}_{i}}^{(tj)}$, we can write:
(4) $G_{{\bf x}_{i}}^{(tj)}=\sum_{r\geq 1}\pi_{r}^{(tj)}({\bf x}_{i})\delta_{\mathbb{\omega}^{(tj)}_{r}},$
@@
for $t=\{0,1\}$ and $j=\{1,\dots,J_{t}\}$, where $\{\pi_{r}^{(tj)}({\bf x}_{i})\}_{r\geq 1}$ and $\{\omega^{(tj)}_{r}\}_{l\geq 1}$ represent infinite sequences of random weights and random kernels’ parameters, respectively. Notably, both random sequences depend on the level of treatment, $t$, and factor, $j$, while the weights also depend on the values ${\bf x}_{i}$ of the confounders.
@@
The elements of the sequence of random parameters $\{\omega_{r}^{(tj)}\}_{r\geq 1}$ are independent and identically distributed, where the $r$-th element of the sequence is the joint parameters of the $r$-th Gaussian distribution in the mixture model in (3), so that $\omega_{r}^{(tj)}=(\eta_{r}^{(tj)},1/\tau_{r}^{(tj)})$. Following the stick-breaking representation (Sethuraman, 1994), the sequence of dependent weights is defined as
$\pi_{r}^{(tj)}({\bf x}_{i})=V^{(tj)}_{r}({\bf x}_{i})\prod_{g<r}\{1-V_{g}^{(tj)}({\bf x}_{i})\},$
where $\{V_{r}^{(tj)}({\bf x}_{i})\}_{r\geq 1}$ are $\{0,1\}$-valued independent stochastic processes.
@@
Among the dependent nonparametric processes, the Dependent Probit Stick-Breaking (DPSB) (Rodriguez and Dunson, 2011) defines the process $\{V_{r}^{(tj)}({\bf x}_{i})\}_{r\geq 1}$ as follows:
$V_{r}^{(tj)}({\bf x}_{i})=\Phi(a_{r}^{(tj)}({\bf x}_{i})),$
(5) $a_{r}^{(tj)}({\bf x}_{i})\sim\mathcal{N}(\alpha_{0r}^{(tj)}+\alpha_{1r}^{(tj)}{\bf x}_{1i}+\dots+\alpha_{pr}^{(tj)}{\bf x}_{pi},1),$

### v3

@@ Abstract.
footnotetext: dafne_\_zorzetto@brown.edu
@@ 1. Introduction
In observational studies, a single cause or exposure can affect multiple outcomes that are often inherently strongly correlated. For instance, in environmental health, air pollution may simultaneously influence several diseases or mortality causes. In genomics, a cancer treatment can alter mutational signature profiles. Similarly, as analyzed in this paper, wildfire smoke represents a specific exposure that can change the chemical composition of ambient air pollutants.
@@
Wildfires have emerged as a critical environmental and public health concern, significantly deteriorating air quality through the emission of smoke and associated pollutants (Childs et al., 2022; Krasovich Southworth et al., 2025). In the United States, wildfire smoke exposure has increased significantly, with the burned area quadrupling over the past four decades, a trend expected to continue as climate conditions become increasingly favorable to fire activity. Wildfire smoke can rise high into the atmosphere and travel long distances, affecting air pollution levels and health risks far from the source (Burke et al., 2021; O’Dell et al., 2021). As a result, wildfire pollution has received considerable attention in environmental health research (Burke et al., 2021; Krasovich Southworth et al., 2025).
@@
A growing body of literature has examined the effects of wildfire smoke on particulate matter (pm2.5), highlighting the need to quantify the proportion of pm2.5 attributable to wildfires (Burke et al., 2021; Childs et al., 2022), given the well-documented adverse health effects of elevated pm2.5 exposure (Carone et al., 2020; U.S. Environmental Protection Agency, 2022). However, wildfire smoke is a complex mixture of pollutants that undergoes physical and chemical transformations during atmospheric transport. Understanding how wildfire smoke influences the concentration levels of various chemical species in pm2.5is therefore a pressing challenge.
@@
While previous studies have focused on the impact of wildfire smoke on individual chemical components —such as phosphorus and nitrogen (Spencer and Hauer, 1991), lead (Odigie and Flegal, 2014), bromine and other metals (Li et al., 2023b)— the recent work of Krasovich Southworth et al. (2025) has provided a more comprehensive analysis by quantifying the contribution of wildfire smoke to the concentration of 27 different chemical species in pm2.5 using a standard linear regression model. Their analysis focuses primarily on accounting for temporal variation and estimating the contribution of wildfire smoke to each chemical species in pm2.5. However, it does not jointly model chemical species, limiting information sharing across species, and does not adopt a causal inference perspective, restricting its ability to disentangle the causal effect of wildfire smoke.
@@
A causal inference framework is crucially needed to rigorously assess and quantify the effect of wildfire smoke on changes in chemical concentration levels, providing a deeper understanding of the associations highlighted in previous studies. A Bayesian perspective is particularly well suited to this setting, as it provides coherent uncertainty quantification for causal estimands, allows prior information to be incorporated in a principled way, and offers a flexible modeling framework for complex data structures (Li et al., 2023a). Moreover, incorporating the high correlation structure among the different chemical species is critical to capturing deeper insights into the composition of smoke and its broader environmental effects.
@@
Traditional Bayesian factor models have focused primarily on priors for factor loadings (e.g., Bhattacharya and Dunson, 2011; Schiavon et al., 2022), typically assuming standard normal distributions for factor scores. Recent works (Zorzetto et al., 2024c; Bortolato and Canale, 2024) have begun exploring priors on factor scores, since they define the subject-specific contribution of each factor to the variation, making them essential in many applications. However, introducing a causal inference framework in Bayesian factor models poses new challenges due to the fundamental missing data problem of causal inference (Rubin, 1980). Our proposed prior addresses this by allowing a covariate-dependent structure in treatment-specific scores, enabling better estimation of latent confounding patterns. In particular, our prior leverages the flexibility of the Dependent Dirichlet Process (DDP) (Quintana et al., 2022), building on a growing body of work in causal inference that employs DDP priors (e.g., Roy et al., 2024; Hu et al., 2023; Zorzetto et al., 2024a), though primarily in the context of scalar outcomes.
A key strength of this model is its flexibility and broad applicability to a wide range of real-world datasets where a treatment or exposure affects multivariate outcomes composed of correlated components. Beyond the wildfire application, the model can be used, for example, to estimate the effects of pm2.5 exposure on cause-specific mortality or to investigate how cancer drugs affect transcriptional profiles.
The paper is organized as follows. Section 2 introduces the causal inference setup, including the estimands of interest, causal assumptions, and identification strategy. Section 3 introduces our Bayesian factor model for causal inference, highlighting the innovation in our prior specification for factor scores and its critical role in this context. In Section 4, we assess the performance of the model performance through simulations. In Section 5, we apply our approach to estimate the causal effect of the presence of wildfire smoke on the concentration of 27 chemical species in pm2.5 across the United States. We conclude the paper with a discussion and future directions in Section 6.
@@ 2. Causal setup
Let $i$ denote the study unit, with $i=1,\dots,n$. For each unit, let $T_{i}\in\{0,1\}$ the binary treatment random variable with observed value $t_{i}$, let ${\bf X}_{i}=(X_{1i},\dots,X_{pi})^{T}$ be the vector of $p$ observed covariates (confounders), which may be either continuous or discrete, and let $\mathbf{Y}_{i}\in\mathbb{R}^{q}$ be the multivariate outcome of dimension $q$. The outcome components are potentially highly correlated, a structure acknowledged by our Bayesian causal factor model. Throughout the paper, bold letters denote vectors.
Following the definition of the Rubin Causal Model (Rubin, 1974), we make the Stable Unit Treatment Value Assumption (SUTVA; Rubin, 1980), which is a combination of two assumptions: no interference among the units, i.e. the potential outcome values from unit $i$ do not depend on the treatment applied to other units, and consistency, i.e. each observed outcome is the individual’s potential outcome under their treatment. Under SUTVA, we define the potential outcomes, for unit $i$, as $\{{\bf Y}_{i}(0),{\bf Y}_{i}(1)\}$, where ${\bf Y}_{i}(t)=(Y_{i1}(t),\dots,Y_{iq}(t))^{T}$ for $t\in\{0,1\}$. Specifically, ${\bf Y}_{i}(0)\in{\mathds{R}}^{q}$ is the $q$-variate outcome when the unit $i$ is assigned to the control group and ${\bf Y}_{i}(1)\in{\mathds{R}}^{q}$ is the $q$-variate outcome when the unit $i$ is assigned to the treatment group. Under consistency, therefore the observed outcome can be written as ${\bf Y}_{i}=(1-T_{i})\,{\bf Y}_{i}(0)+T_{i}\,{\bf Y}_{i}(1).$ In practice, for $i=1,\dots,n$, we observe only ${\bf y}_{i}\in{\mathds{R}}^{q}$, the realization of the random variable ${\bf Y}_{i}=(Y_{i1},\dots,Y_{iq})$. Conversely, we do not observe the missing outcome ${\bf Y}_{i}^{mis}\in{\mathds{R}}^{q}$ defined as ${\bf Y}_{i}^{mis}=T_{i}\,{\bf Y}_{i}(0)+(1-T_{i})\,{\bf Y}_{i}(1)$.
To concisely define the correlated nature of the multidimensional outcome, for each unit $i$, we assume an unmeasured variable ${U}_{i}$ that affects the treatment $T_{i}$ and the potential outcome $\{{\bf Y}_{i}(0),{\bf Y}_{i}(1)\}$. Under the assumption of our proposed Bayesian causal factor model, formally introduced in Section 3. Bayesian Causal Factor Model, we model both factual and counterfactual latent factors $\{{\bf L}_{i0},{\bf L}_{i1}\}$ that capture the indirect effect of $U_{i}$ on the potential outcomes $\{{\bf Y}_{i}(0),{\bf Y}_{i}(1)\}$, as illustrated in Figure 1 Figure 1 shows the three versions of the resulting conditional independence structure considered in our model.
@@
In our environmental data application, the latent factors $L_{t}$, identified in our factor model, explain the observed correlation structure among chemical species. While the latent variables denoted by $U$ may represent specific pollution sources such as traffic, metalworking industries, energy production, or soil dust (see e.g. Park et al., 2014), which are often impractical to measure comprehensively, but may determine the latent correlation structure.
@@
Positivity. Each unit $i\in\{1,\dots,n\}$ has a non-zero probability of receiving each treatment level, such that $0<\Pr(T_{i}=1\mid{\bf X}_{i}={\bf x}_{i})<1.$
Conditional ignorability. The assignment of treatment is random in each group of units characterized by observed covariates $\bf x$} and latent factors scores ${\bf l}_{0}$ and ${\bf l}_{1}$, such that $\{{\bf Y}_{i}(1),{\bf Y}_{i}(0)\}\perp T_{i}\mid\{{\bf X}_{i}={\bf x}_{i},{\bf L}_{i0}={\bf l}_{i0},{\bf L}_{i1}={\bf l}_{i1}\},$ for each unit $i\in\{1,\dots,n\}$.
Indirectness of unmeasured confounding effects. The unmeasured variable $U$ affects the potential outcomes only indirectly, through the latent factor and the observed confounders, such as $\mathds{P}({\bf Y}(1),{\bf Y}(0)\mid{\bf X},{\bf L}_{0},{\bf L}_{1},U)=\mathds{P}({\bf Y}(1),{\bf Y}(0)\mid{\bf X},{\bf L}_{0},{\bf L}_{1}).$ This assumption can be visualized in Figure 1 as the absence of a direct arrow between the $U$ and $\bf Y$}.
@@
${\mathds{E}}[{\bf Y}(1)-{\bf Y}(0)]=\int_{\mathcal{X}}\int_{\mathbb{R}^{J_{0}}}\int_{\mathbb{R}^{J_{1}}}\bigg({\mathds{E}}[{\bf Y}\mid T=1,{\bf X}={\bf x},{\bf L}_{0}={\bf l}_{0},{\bf L}_{1}={\bf l}_{1}]$
$\quad\quad-{\mathds{E}}[{\bf Y}\mid T=0,{\bf X}={\bf x},{\bf L}_{0}={\bf l}_{0},{\bf L}_{1}={\bf l}_{1}]\bigg)\times\mathds{P}({\bf L}_{0}={\bf l}_{0},{\bf L}_{1}={\bf l}_{1}\mid{\bf X}={\bf x})\mathds{P}({\bf X}={\bf x})d{\bf l}_{0},d{\bf l}_{1}d_{\bf x};$
@@
${\mathds{E}}[{\bf Y}\mid t,x,l_{0},l_{1}]=\int_{\mathcal{U}}{\mathds{E}}[{\bf Y}\mid t,x,l_{0},l_{1},u]\mathds{P}(u\mid t,x,l_{0},l_{1})du$
(1) $\quad\quad\quad\quad\quad\quad\quad\;=\int_{\mathcal{U}}\frac{{\mathds{E}}[{\bf Y}\mid t,x,l_{0},l_{1}]\mathds{P}(l_{0},l_{1}\mid u,x)\mathds{P}(t\mid x,u)\mathds{P}(x,u)}{\int_{u}\mathds{P}(l_{0},l_{1}\mid u,x)\mathds{P}(t\mid x,u)\mathds{P}(x,u)du}du;$
@@ 3. Bayesian Causal Factor Model
To address the multivariate outcome, we propose an approach that adopt a factor regression model (Carvalho et al., 2008), which uses factor analysis to model the potential outcomes given the treatment and the confounders. Rather than performing multiple univariate analyses, our factor analysis approach (i) improves statistical power by leveraging the correlation structure in the outcome, and (ii) facilitates the recovery of information from the unmeasured variable U{U}, under the three cases considered in Figure 1.
@@
(2) $g_{t}({\bf x}_{i},{\bf l}_{it})=\boldsymbol{\mu}_{t}+{\bf B}_{t}{\bf x}_{i}+\Lambda_{t}{\bf l}_{it}+\boldsymbol{\xi}_{it},\quad\boldsymbol{\xi}_{it}\sim\mathcal{N}_{q}(0,\Psi_{t}).$
The outcome ${\bf Y}_{i}$ depends on treatment $t$, covariates ${\bf x}_{i}$, and treatment-specific latent factors ${\bf l}_{it}\in\mathbb{R}^{J_{t}}$, where $J_{t}$ denotes the number of factors for treatment $t$ (assumed unknown). The matrix $\Lambda_{t}\in\mathbb{R}^{q\times J_{t}}$ represents treatment-specific factor loadings, where the element $\lambda_{tjh}$ denotes the entry in row $j$ and column $h$, and $\Psi_{t}\in\mathbb{R}^{q\times q}$ is assumed to be a diagonal matrix with elements $(\psi_{1t},\dots,\psi_{qt})$ as in standard factor analysis literature (Thurstone, 1931; Anderson-Cook, 2004).
@@
where $\boldsymbol{\beta}_{tj}$ denotes the $j$-th row of the matrix ${\bf B}_{t}$ in (2), such that $\boldsymbol{\beta}_{tj}=\{\beta_{tj1},\dots,\beta_{tjp}\}$ and $\beta_{tjk}$ is the element $({\bf B}_{t})_{j,k}$ with $j\in\{1,\dots,q\}$ and $k\in\{1,\dots,p\}$. We assume that $\Sigma_{m_{t}}$ and $\Sigma_{\beta_{t}}$ are diagonal matrices for each $t$.
@@
The flexibility of the DDP has made it widely adopted in causal inference settings involving scalar outcomes. For example, Roy et al. (2024) leverage the enriched Dirichlet process to estimate causal effects for mediation analysis, while Zorzetto et al. (2024a) specify a DDP prior for heterogeneous treatment effect.
In our factor model, our DDP prior has two characteristics: (i) it can written as infinite mixture of Gaussian distributions, and (ii) the probability of assignment to each mixture component depends on the covariates XX—called cluster allocation in the DDP literature. Moreover, we specify a DDP prior that allows us to capture heterogeneity at both the factor and treatment levels. In fact, treatment-specific factor scores quantify the score/importance that each unit expresses on the corresponding factor under the specific treatment level. In the environmental application discussed in Section 5. Environmental Application, each factor score of a specific chemical indicates its contribution to explain the correlation among chemicals within the corresponding latent factor (reflecting underlying sources of metal and organic emissions). Notably, each factor score can vary depending on the presence or absence of wildfire smoke, reflecting how wildfire exposure alters the chemical composition of the air.
@@
(3) $\{l_{it{\color[rgb]{0,0,0}h}}\mid{\bf x}_{i},t\}\sim\int_{\Omega}{\mathcal{H}}(\cdot;\omega)dG_{{\bf x}_{i}}^{(t{\color[rgb]{0,0,0}h})}(\omega),\quad G_{{\bf x}_{i}}^{(t{\color[rgb]{0,0,0}h})}\sim\Pi_{{\bf x}_{i}}^{(t{\color[rgb]{0,0,0}h})},$
for ${\color[rgb]{0,0,0}h}\in\{1,\dots,J_{t}\}$ of units $i\in\{1,\dots.n\}$, for each $t\in\{0,1\}$. The random probability measure $G_{{\bf x}_{i}}^{(t{\color[rgb]{0,0,0}h})}$ depends on confounders ${\bf x}_{i}$ and is specific to a treatment level $t$ and a factor $\color[rgb]{0,0,0}h$}. Assuming its prior $\Pi_{{\bf x}_{i}}^{(t{\color[rgb]{0,0,0}h})}$ where $\Pi_{{\bf x}_{i}}^{(t{\color[rgb]{0,0,0}h})}$ is a DDP (Quintana et al., 2022), we define a dependent infinite mixture distribution as the probability density distribution of the factors.
@@
Following a single-atom DDP (Quintana et al., 2022) characterization of the random measure $G_{{\bf x}_{i}}^{(t{\color[rgb]{0,0,0}h})}$, we can write:
(4) $G_{{\bf x}_{i}}^{(t{\color[rgb]{0,0,0}h})}=\sum_{r\geq 1}\pi_{r}^{(t{\color[rgb]{0,0,0}h})}({\bf x}_{i})\delta_{\mathbb{\omega}^{(t{\color[rgb]{0,0,0}h})}_{r}}$,
@@
for $t=\{0,1\}$ and ${\color[rgb]{0,0,0}h}=\{1,\dots,J_{t}\}$, where $\{\pi_{r}^{(t{\color[rgb]{0,0,0}h})}({\bf x}_{i})\}_{r\geq 1}$ and $\{\omega^{(t{\color[rgb]{0,0,0}h})}_{r}\}_{l\geq 1}$ represent infinite sequences of random weights and random kernels’ parameters, respectively. Notably, both random sequences depend on the level of treatment, $t$, and factor, $\color[rgb]{0,0,0}h$}, while the weights also depend on the values ${\bf x}_{i}$ of the confounders.
@@
The elements of the sequence of random parameters $\{\omega_{r}^{(t{\color[rgb]{0,0,0}h})}\}_{r\geq 1}$ are independent and identically distributed, where the $r$-th element of the sequence is the joint parameters of the $r$-th Gaussian distribution in the mixture model in (3), so that $\omega_{r}^{(t{\color[rgb]{0,0,0}h})}=(\eta_{r}^{(t{\color[rgb]{0,0,0}h})},1/\tau_{r}^{(t{\color[rgb]{0,0,0}h})})$. Following the stick-breaking representation (Sethuraman, 1994), the sequence of dependent weights is defined as
$\pi_{r}^{(t{\color[rgb]{0,0,0}h})}({\bf x}_{i})=V^{(t{\color[rgb]{0,0,0}h})}_{r}({\bf x}_{i})\prod_{g<r}\{1-V_{g}^{(t{\color[rgb]{0,0,0}h})}({\bf x}_{i})\},$
where $\{V_{r}^{(t{\color[rgb]{0,0,0}h})}({\bf x}_{i})\}_{r\geq 1}$ are $\{0,1\}$-valued independent stochastic processes.
@@
Among the dependent nonparametric processes, the Dependent Probit Stick-Breaking (DPSB) (Rodriguez and Dunson, 2011) defines the process $\{V_{r}^{(t{\color[rgb]{0,0,0}h})}({\bf x}_{i})\}_{r\geq 1}$ as follows:
(5) $V_{r}^{(t{\color[rgb]{0,0,0}h})}({\bf x}_{i})=\Phi(a_{r}^{(t{\color[rgb]{0,0,0}h})}({\bf x}_{i})),\quad a_{r}^{(t{\color[rgb]{0,0,0}h})}({\bf x}_{i})\sim\mathcal{N}(\alpha_{0r}^{(t{\color[rgb]{0,0,0}h})}+\alpha_{1r}^{(t{\color[rgb]{0,0,0}h})}{\bf x}_{1i}+\dots+\alpha_{pr}^{(t{\color[rgb]{0,0,0}h})}{\bf x}_{pi},1),$